A car travels 250 kilometrs in the same time that a truck travels 225 kilometers. If the speed of the car is 8 kilometers per hour faster than the speed of the truck, find both speeds.
What is the equation?

Answer by mananth(16946)

Truck----x------225 km
Car ----x+8 ----- 250 km

t=d/r
225/x = 250 /(x+8)
Cross multiply
225(x+8 ) = 250 * x
225 x + 1800 = 250 x
250 x -225 x = 1800
25 x = 1800
/ 25
x = 72 Truck kph
80 Car kph